OpenStudy (anonymous):

This formula converts Fahrenheit to Celsius. C = 5/9 (F−32) Which is a solution for F?

OpenStudy (anonymous):

9C/5 +32 =F

OpenStudy (anonymous):

C = 5/9 (F-32) F - 32 = 9/5 C F = (9/5 C) + 32

OpenStudy (anonymous):

F=(C(9/5)) + 32

OpenStudy (anonymous):

\[f=32+C(\frac{9}{5})\]